OpenStudy (anonymous):

Evaluate. (5 • 8 – 12)(36 ÷ 9 • 4)

OpenStudy (anonymous):

lets work on the first part (5*8-12)

OpenStudy (anonymous):

5 times 8 = 40 - 12 =28

OpenStudy (anonymous):

the next one (36 divided by 9*4)

OpenStudy (anonymous):

36 didived by 9 = 4*4 = 16

OpenStudy (anonymous):

so 16+28=44

OpenStudy (anonymous):

so ur answer is 44

OpenStudy (anonymous):

i tohught it was multply next

OpenStudy (anonymous):

oh i mean 16*28=.....

OpenStudy (anonymous):

392

OpenStudy (anonymous):

448 is what i got

OpenStudy (anonymous):

i got it now i did a different number

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

your welcome